OpenStudy (buttermequeasy):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (buttermequeasy):

Help please?

OpenStudy (mertsj):

Plug in the point (-3,-1) and see which sentence is true.

OpenStudy (mertsj):

Which one is it?

OpenStudy (buttermequeasy):

hold on

OpenStudy (buttermequeasy):

The third one? y = 4x + 7

OpenStudy (mertsj):

-1=4(-3)+7 -1=-12+7 -1=-5 Is that true?

OpenStudy (mertsj):

It's much easier to help someone if they actually answer you.

OpenStudy (buttermequeasy):

Sorry.

OpenStudy (buttermequeasy):

I was doing it backwards.

OpenStudy (buttermequeasy):

y = 4x + 11 The second one?